suppose tacos cost $4 each and hot dogs cost $2. Moreover, this mexican restaurant has a special deal. after you've bought 6 burritos, you can each extra burrito for $1.50. Assume that tacos are on the vertical axis and the burritos on the horizontal axis. what is the marginal rate of transformation when 8 burritos are purchased?
Answer #1

PT = 4, PB= 2

If B>6, then PB` = 1.5

B on X axis, then MRTS = PB/PT

.

if B = 8, MRTS = PB`/PT

MRTS = 1.5/4 = .375